### v1

# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Medical Image

Abstract
Semantic medical image segmentation is a crucial part of both scientific research and clinical care. With enough labelled data, deep learning models can be trained to accurately automate specific medical image segmentation tasks. However, manually segmenting images to create training data is highly labor intensive. In this paper, we present ScribblePrompt, an interactive segmentation framework for medical imaging that enables human annotators to segment unseen structures using scribbles, clicks, and bounding boxes. Scribbles are an intuitive and effective form of user interaction for complex tasks, however most existing methods focus on click-based interactions. We introduce algorithms for simulating realistic scribbles that enable training models that are amenable to multiple types of interaction. To achieve generalization to new tasks, we train on a diverse collection of 65 open-access biomedical datasets – using both real and synthetic labels. We test ScribblePrompt on multiple network architectures and unseen datasets, and demonstrate that it can be used in real-time on a single CPU. We evaluate ScribblePrompt using manually-collected scribbles, simulated interactions, and a user study. ScribblePrompt outperforms existing methods in all our evaluations. In the user study, ScribblePrompt reduced annotation time by 28% while improving Dice by 15% compared to existing methods. We showcase ScribblePrompt in an online demo and provide code at scribbleprompt.csail.mit.edu.
Figure 1: Example predictions with iterative simulated interactions (steps). ScribblePrompt enables iterative interactive segmentation of unseen tasks using scribbles, clicks, and bounding boxes. We show representative examples with simulated interactions that reflect the diversity of tasks in our evaluation data. At each step, we visualize positive scribble and click inputs in green, negative scribble and click inputs in red, bounding box inputs in yellow, and the predicted segmentation in blue.
1 Introduction
Semantic segmentation of medical images is an essential step in a wide range of biomedical research and clinical care pipelines. Deep learning has become the predominant method to automate existing segmentation tasks [108, 53, 123, 94]. However, training accurate models is a large undertaking that depends on having diverse images carefully annotated by skilled experts. Biomedical researchers and clinicians often target novel segmentation tasks, requiring the delineation of either new regions of interest [142] or new imaging modalities [113, 91].
Most medical image segmentation systems used in practice provide simplistic editing tools, sometimes with intensity-based algorithmic assistance, such as ITK-SNAP [141]. Recent research has focused on learning-based interactive segmentation systems in which the annotator iteratively collaborates with a model to complete the segmentation, providing prompts, such as clicks, and receiving predicted segmentation(s) that they can correct with additional prompting.
Given the diversity of biomedical imaging available, interactive segmentation tools would be particularly useful when able to segment previously unsegmented biomedical domains, regions of interest, and image types. However, existing methods in this space have not been widely adopted in practice, mostly because they are useful only in limited circumstances or for specific segmentation targets such as abdominal organs in CT [80, 111, 131, 100, 130, 150]. Recent foundation models tackle more general purpose semantic segmentation [61, 151, 132, 133], but also do not generalize to new medical imaging tasks without fine-tuning for the desired tasks [90, 50, 44, 117, 82, 25, 135]. Other methods avoid fine-tuning but require extensive “prompts” involving already labeled examples [20, 28].
Most prior work on interactive segmentation has been limited to click-based [120, 76, 23, 13, 80, 61, 137, 70, 34] and bounding box-based interactions [138, 61, 106], often requiring users to follow specialized click protocols [144] such as clicking extreme points [85] or interior margin points [80]. Scribbles are a natural mode of interaction; a large, rough, scribble is often easier and faster to perform than a series of carefully placed clicks. Scribbles provide more information than clicks, which is helpful for segmenting complex objects, with minimal increase in user effort. Unfortunately, few existing methods employ scribbles, because of the challenges of acquiring or simulating realistic scribbles for training and evaluation [24].
In this paper, we present ScribblePrompt, an interactive segmentation framework that:
  • enables scribble inputs (along with clicks and bounding boxes), to give users more flexibility. To train the model, we devise ways to simulate diverse and realistic scribbles.
  • enables generalization to unseen segmentation tasks and medical imaging datasets. In addition to training on a large collection of datasets, we propose augmentation strategies using synthetic labels.
  • facilitates interactive refinement of predictions. We simulate iterative interactions during training, and train a single network to handle both the initial prediction and refinement of predictions based on user corrections.
  • is highly efficient, enabling real-time inference, even on a single CPU.
We demonstrate that ScribblePrompt is more accurate than existing methods when tested on manually-collected scribbles and simulated interactions (Figure 1) on twelve datasets. Through a user study with experienced annotators, we show that ScribblePrompt’s ability to segment accurately based on bounding boxes, clicks and scribbles provides valuable flexibility to users, reducing annotation time by 28% while increasing Dice by 15% compared to previous methods.
We release an interactive tool, code, and model weights at scribbleprompt.csail.mit.edu. We will also release new multi-annotator segmentation labels created using ScribblePrompt for select datasets.
2 Related Works
Interactive Medical Image Segmentation. Deep learning-based techniques for interactive segmentation of medical images [80, 100, 130, 131] have gained popularity in research over traditional intensity-based methods [141, 40, 17, 27, 126]. Many existing methods focus on using human interaction to improve the accuracy of segmentation for tasks where some labelled data exists, but fully-automatic segmentation methods fail to produce accurate segmentations [150, 131, 129, 80]. These methods assume that the model will be used to perform the same segmentation task(s) it was trained on, leading to domain-specific solutions [131, 7, 75, 96, 106, 129]. A few methods generalize to new, but similar, classes [111] and modalities [80, 130] to those seen in training. In contrast, we design ScribblePrompt to help annotators segment a wide range of medical images at inference time.
Generalist Models. Recent vision foundation models employ prompting to reduce the burden of manual segmentation and enable generalization to new tasks. These models, trained on large collections of natural image data, segment potentially-unseen structures specified by spatial prompts [61, 151], text descriptions [104, 61, 151], or examples [132, 133, 20, 28]. Some of these models [61, 151] can be used for interactive segmentation, however they are optimized for performance on natural imaging tasks.
Fine-tuning models initially developed using natural images is often unhelpful for specific tasks in the medical imaging domain [105]. The Segment Anything Model [61], a foundation model trained on natural images, performs well at medical image segmentation tasks with clear boundaries (e.g., organ segmentation in abdominal CT), but performs poorly on tasks and modalities involving more subtle boundaries (e.g., deep structures in Brain MRI) [50, 90, 117, 44]. Several recent or concurrent papers fine-tune SAM for specific tasks or modalities [97, 73, 128, 58, 49, 143]. A few fine-tune SAM to segment medical images from multiple-modalities with bounding boxes [82], clicks [135] or both [25]. However, as we show in our experiments, these models don’t perform well on many unseen tasks, including those with thin structure (e.g., blood vessels).
Several natural imaging methods simulate iterative user interactions and condition the model on the previous predictions to train a single network to make the initial prediction and perform refinement [120, 61, 119, 18, 34, 76]. This approach is appropriate for natural images where there is ample data and typically one modality. However, such methods have not previously been developed for the medical imaging domain, which is more fragmented because of specialized regions of interest and different modalities. We assemble a large diverse collection of medical imaging datasets and develop models that facilitate iterative interactive segmentation of unseen medical imaging tasks.
User Interaction. Users can prompt interactive segmentation systems in many different ways, such as bounding boxes [110, 139, 61, 13, 130, 106], scribbles [131, 24, 129], and clicks [131, 80, 13, 76, 120]. Few interactive segmentation models incorporate multiple types of inputs. Some works have explored training with more specialized types of clicks [144], such as extreme points [85, 109], interior margin points [80] and center clicks [120, 137] to maximize the information per interaction. However, specialized interactions often require more user time, and models trained only with such interactions are less robust to deviations from the interaction protocol [139, 120, 13]. In contrast, we develop a user-friendly model that performs well under a variety of interaction scenarios.
Scribbles are an intuitive form of interaction, however few interactive segmentation works utilize them, because simulating realistic scribbles is challenging. Many intensity-based image segmentation methods that do not require training can predict segmentations from scribbles, however they require extensive interactions to produce accurate segmentations [40, 17]. Previous deep-learning models trained for scribble-based segmentation either relied on collecting large datasets of manually drawn scribbles [150, 71] or simplistic approaches such as sampling random points [9, 6, 131] or Bezier curves [3]. One work explored simulating more complex scribbles for natural image segmentation [24]. We build on these concepts to create a new scribble simulation engine for training and evaluating scribble-based segmentation methods.
3 ScribblePrompt Method
We present an interactive segmentation method that can generalize to new medical modalities and regions of interest. We describe the problem formulation, and present the important aspects of the ScribblePrompt framework: (i) simulation of interactions during training, (ii) augmentation with synthetic labels during training to encourage generalization, and (iii) an efficient architecture for real-time interaction.
3.1 Problem Formulation
Let $t$ be a segmentation task consisting of an image and segmentation pair, $\{x^{t},y^{t}\}$. At step $i$, given an image $x^{t}$, a set of user interactions $u_{i}$, and previous prediction $\hat{y}^{t}_{i-1}$, we learn function $f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})$ that produces a segmentation $\hat{y}_{i}$ and has learnable parameters $\theta$. The set of interactions $u_{i}$, which may include positive or negative scribbles, positive or negative clicks, and bounding boxes, is provided by a user who has access to the previous prediction, $\hat{y}_{i-1}^{t}$.
We minimize the difference between $y^{t}$ and the series of iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$,
$\small\mathcal{L}(\theta;\tau)=\mathbb{E}_{t\in\tau}\left[\mathbb{E}_{(x^{t},y^{t})\in t}\left[\sum_{i=1}^{k}\mathcal{L}_{Seg}\left(y^{t},f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})\right)\right]\right]\,$ (1)
where $\mathcal{L}_{Seg}$ is a supervised segmentation loss.
During training, we sample a task $t$ from training data collection $\tau$, an image $x^{t}$, and segmentation map $y^{t}$. We simulate a possible set of interactions $u_{i}$ based on $y^{t}$, and predict $\hat{y}_{i}^{t}=f_{\theta}(x^{t},u_{i},\hat{y}^{t}_{i-1})$. We simulate the next set of interactions $u_{i+1}$ based on the error region $\varepsilon_{i}^{t}=y^{t}-\hat{y}_{i}^{t}$, and repeat for $k$ iterations. In the following sections, we describe strategies for simulating $u_{i}$, collecting $\tau$, sampling $(x^{t},y^{t})$, building $f_{\theta}$ and optimizing $\mathcal{L}(\theta;\tau)$.
Figure 2: Simulated scribbles and clicks. Positive interactions (green) are simulated on the segmentation label $y^{t}$ (blue), while negative interactions (red) are simulated on the background $1-y^{t}$.
3.2 Prompt Simulation
To develop robustness to user interactions $u_{i}$, we introduce algorithms for simulating scribbles, clicks, and bounding box inputs during training. Each scribble and click strategy (Figure 2) can be applied to a ground truth segmentation label $y^{t}$ to simulate positive interactions, or to the background $1-y^{t}$ to simulate negative interactions. We simulate positive and negative correction scribbles or clicks by applying the same strategies to the false negative region $\varepsilon_{i}^{t}>0$ and false positive region $\varepsilon_{i}^{t}<0$ for error region $\varepsilon_{i}^{t}$.
Scribbles. Given binary mask $y\in\{0,1\}^{h\times w}$, we simulate a scribble mask $s\in[0,1]^{h\times w}$ by first generating clean scribbles, and then corrupting them. We illustrate this process in Appendix B.1.1. We start with either:
  1. (i)
    Line Scribbles: We draw random lines by sampling two end points from $\{(u,v)|y_{uv}=1\}$.
  2. (ii)
    Centerline Scribbles: We simulate scribbles drawn in the center of the label. We use a thinning algorithm [145] to draw a 1-pixel wide skeleton of $y$.
  3. (iii)
    Contour Scribbles: We simulate a rough contour of the desired segmentation within the boundaries of the mask. We first blur the reverse mask $\overline{y}=1-y$ to reduce the size of the label such that $\tilde{y}:=1-\max(\overline{y},\overline{y}\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{y}<h$ sampled in some intensity range $h\sim U[\tilde{y}_{min},\tilde{y}_{max}]$ and use an edge detection algorithm to extract a contour inside the boundary of the mask.
To limit the size of the centerline and contour scribbles, we apply a random mask that breaks the scribbles into smaller parts. We generate the random mask by sampling a smooth noise image $p$, where each pixel is sampled independently from $\mathcal{N}(\mu_{p},\sigma_{p}^{2})$, applying Gaussian blur, and then thresholding it at $\mu_{p}$.
Then, we warp the scribble mask $s$ using a random deformation field $\phi$ to vary the scribble shape and thickness. We ensure the resulting scribble is consistent with mask $y$, by multiplying the warped scribble mask $s\circ\phi$ by $y$.
Clicks. Given a binary mask $y\in\{0,1\}^{h\times w}$, we simulate $n\sim U[n_{min},n_{max}]$ clicks at a time using one of three strategies illustrated in Figure 2:
  1. (i)
    Random clicks: We randomly sample clicks from all pixels in the given region $\{(u,v)|y_{uv}=1\}$.
  2. (ii)
    Center clicks: We place clicks at the center of each component of the label. First, we create a multi-label mask $m\in\{1,\dots,C\}^{h\times w}$ identifying the $C$ disconnected components of label mask $y$. We then sample pixel coordinates from the set of center points of the components based on the euclidean distance transform, similar to [120, 137].
  3. (iii)
    Interior border region clicks: We sample clicks from a border region inside the boundary of the mask. We first blur the reverse mask $\overline{y}=1-y$ to reduce the size of the label such that $\tilde{y}:=1-\max(\overline{y},\overline{y}\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel. We then sample click coordinates from $\{(u,v)|\tilde{y}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{y}_{min},\tilde{y}_{max})$ are thresholds sampled in some intensity range. We illustrate this process in Appendix B.1.2.
Since users are inclined to spread out their clicks, we impose a minimum separation of a few pixels between random clicks and border region clicks sampled in the same step.
Bounding Boxes. We compute the minimum bounding box that encloses the label $y^{t}$, and enlarge each dimension by $r\sim U[0,20]$ pixels to account for human error.
Iterative Training. During the first step ($i=1$) we sample the combination of interactions and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The Initial set of interactions $u_{1}$ are simulated based on the ground truth label $y^{t}$. In subsequent steps, correction scribbles or clicks are sampled from the error region between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth segmentation $y^{t}$. Since a user can make multiple corrections before refreshing the prediction, during training we sample $n_{cor}\sim U[n_{min},n_{max}]$ correction prompts (scribbles or clicks) per step.
3.3 Data
We build on MegaMedical [20] to construct a collection of 77 open-access medical imaging datasets for training and evaluation. The collection includes a diverse array of biomedical domains, such as eyes [48, 69, 83, 102, 122], thorax [112, 116, 118, 114, 103], spine vertebrae [148, 78, 114, 134], cells [149, 77, 21, 36, 22, 35], skin [26], abdominal [15, 16, 45, 54, 57, 65, 66, 68, 74, 79, 103, 118, 84, 121, 107], neck [62, 101, 95, 99], brain [10, 37, 46, 63, 64, 86, 87, 89, 118, 4], bones [114, 42, 134], teeth [1, 51] and lesions [5, 146, 147, 118].
We define a 2D segmentation task as a combination of dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice for each volume.
Task Diversity. During training we sample image, segmentation pairs hierarchically i.e., by dataset and modality, axis and then label, to balance training on datasets of different sizes. To increase the diversity of segmentation tasks, we apply data augmentation to both the input image and sampled segmentation prior to simulating the user interactions.
Synthetic Labels. To limit overfitting to specific segmentation tasks, we introduce synthetic labels. During training, for a given sample $(x^{t},y^{t})$, with probability $p_{synth}$ we replace $y^{t}$ with a synthetic label $y^{\prime}$. Given image $x^{t}$ we generate a synthetic label $y^{\prime}$ by applying a superpixel algorithm [33] with scale parameter $\lambda\sim U[1,\lambda_{max}]$ to partition the image into a multi-label mask of $k$ superpixels $z\in\{1,\dots,k\}^{n\times n}$ and then randomly selected a superpixel $y^{\prime}:=\mathbbm{1}(z=k)$.
3.4 Network
We demonstrate ScribblePrompt using an efficient fully-convolutional architecture similar to a Unet [108], ScribblePrompt-Unet, and also a vision transformer-based architecture, the Segment Anything Model (SAM) [61], ScribblePrompt-SAM. We provide code and weights for both models at scribbleprompt.csail.mit.edu.
ScribblePrompt-Unet. We use an 8 layer CNN following a decoder-encoder structure similar to the popular Unet architecture [108] without Batch Norm. Each convolutional layer has 192 features and uses PReLu activation [43].
ScribblePrompt-SAM. For ScribblePrompt-SAM we fine-tune the smallest SAM model (ViT-b). We use ViT-b because we observed the smallest backbone (ViT-b) had similar performance to the larger backbones on the validation set. We freeze all weights except for the decoder.
The SAM architecture takes bounding box and click inputs as lists of ${(x,y)}$ coordinates, and the low-resolution logits of the previous prediction $y_{i-1}$. To adapt SAM for scribbles, we represent positive/negative scribbles as a set of positive/negative clicks on every pixel of the scribbles.
Loss. We minimize eq. (1) where $\mathcal{L}_{Seg}$ is a linear combination of soft Dice Loss [32] and Focal Loss [72]. In preliminary experiments (Appendix B.2), training with Soft Dice Loss [32] alone or a linear combination of Soft Dice Loss and Binary-Cross Entropy Loss resulted in slightly lower Dice scores.
4 Experimental Setup
We compare ScribblePrompt-Unet and ScribblePrompt-SAM to previous interactive segmentation methods with both a user study with experienced annotators, and with manual and simulated interactions.
Data. We use 65 datasets during training. We partition another nine datasets, ACDC [14], BUID [5], BTCV Cervix [66], DRIVE [122], HipXRay [42], PanDental [1], SCD [103], SpineWeb [148], and WBC [149], into a validation set (used in model selection, but not training) and a test set (used only for final evaluation). Each dataset was split into 60% train, 20% validation, and 20% test by subject. Three additional datasets, TotalSegmentator [134], SCR [125], and COBRE [4], were used only for final evaluation. We selected these datasets to cover a variety of modalities (MRI, CT, Ultrasound, OCT, Microscopy) and anatomical regions (Brain, Teeth, Bones, Organs, Muscles, Heart, Thorax, Cells), including both healthy anatomy and lesions. We report results on twelve sets of data covering 608 tasks and eight modalities.
We resized all images to 128x128 and re-scaled intensities to [0,1]. For methods using the SAM architecture, we convert the images to RGB and apply the pixel normalization scheme in [61]. We train with $p_{synth}=0.5$ and apply the same data augmentations as [20]. We provide more details on the data in Appendix F.
Training. We simulate five iterations of interactions per example and set the maximum number of interactions per iteration to three. We set the minimum number of initial negative prompts to zero, and the minimum initial positive and correction prompts to one. During the first iteration, we sample the combination of prompt types (clicks, scribbles, bounding boxes) and then sample the number of positive and negative interactions for each type. In each subsequent iteration, we simulate either scribble corrections or click corrections with equal probability. We selected these values based on what we believe is a reasonable number of steps and interactions a user might perform.
We use the Adam optimizer [59] and train with a learning rate of $0.0001$ until convergence. We use a batch size of 8 for ScribblePrompt-Unet. For ScribblePrompt-SAM we use a batch size of 1, because of memory constraints.
Figure 3: Example predictions. Left to right: SAM, SAM-Med2d, MIDeepSeg, ScribblePrompt-Unet, and ScribblePrompt-SAM. Top: predictions after one step of manual scribbles. Bottom: predictions after five steps of simulated interactions (one center click or bounding box, followed by one center correction click per step). Representative examples were selected to show the diversity of tasks in our evaluation.
4.1 Baselines
We compare to the following existing methods for iterative interactive segmentation of medical images.
SAM [61]: We evaluate the smallest (ViT-b) and largest (ViT-h) versions of the Segment Anything Model (SAM), trained on a large natural image dataset. SAM takes bounding boxes and clicks as input.
SAM-Med2d [25]: SAM-Med2d is a SAM model with ViT-b backbone and additional adapter layers in the image encoder. SAM-Med2d was fine-tuned on a collection of public and private datasets containing 4.6M medical images and 19.7M segmentations [^1] SAM-Med2d used three of our evaluation datasets (ACDC, BTCV and TotalSegmentator) as training datasets [140]., using bounding boxes and iterative clicks as input. Following [25], we evaluate SAM-Med2d both with and without adapter layers.
MIDeepSeg [80]: MIDeepSeg is an interactive segmentation framework designed to generalize to unseen tasks on medical images. MIDeepSeg takes interior margin points (positive clicks) as initial inputs. It crops the image based on those points and uses a CNN to make an initial prediction. The prediction can be refined with additional clicks using conditional random fields. We evaluate the pre-trained model, which was developed on Placenta T2 MRI.
5 Experiments
We evaluate all methods using both manual and simulated interactions. Simulated interactions enable us to test on a large volume of images and tasks. However, simulations do not always match user behavior. We address this by (1) collecting a small dataset of manual scribbles for validation and (2) conducting an interactive user study.
5.1 Manual Scribbles
Setup. To validate the proposed scribble simulation procedure, we collected a dataset of manual scribbles from three annotators for seven segmentation tasks with a total of 93 examples. The annotators were shown five training examples with the ground truth segmentation per task and instructed to draw positive and negative scribbles to indicate the region of interest on new images.
In this evaluation, we make a prediction given a set of positive and negative scribbles, simulating a non-iterative scenario where the annotator draws several scribbles before running inference. We evaluate accuracy using Dice score [32]. MIDeepSeg’s initial prediction network only takes positive inputs; we report results on both MIDeepseg’s initial prediction from the positive scribbles, and after applying MIDeepSeg’s refinement procedure using the negative scribbles as if they were corrections.
Results. ScribblePrompt-SAM and ScribblePrompt-Unet produce the most accurate segmentations in a single step from the manual scribbles, with an average Dice of 0.87 and 0.81, respectively (Table 1). Scribbles are particularly helpful for large segmentations covering areas with varying intensity, such as the jaw in dental X-Rays (Figure 3).
SAM and SAM-Med2d do not generalize well to scribble inputs (which they were not trained for). The initial predictions from MIDeepSeg’s network are poor (0.68 Dice), but improve to 0.81 Dice after applying the refinement procedure, suggesting that for some examples in our manual scribble dataset non-DL methods may be sufficient. We evaluate the generalizability of all methods on more tasks through experiments with simulated interactions.
Table 1: Evaluation on manual scribbles. Mean Dice of segmentations ($n=93$) predicted from manual scribbles on examples from validation datasets. The ScribblePrompt models, which were trained with simulated scribbles (as well as clicks and boxes), have the highest Dice from a single prediction.
Model Dice (mean $\pm$ ci)
SAM (ViT-b) $0.40\pm 0.05$
SAM (ViT-h) $0.57\pm 0.05$
SAM-Med2d w.o. adapter $0.52\pm 0.05$
SAM-Med2d w/ adapter $0.30\pm 0.06$
MIDeepSeg $0.68\pm 0.04$
MIDeepSeg (with refinement) $0.81\pm 0.03$
ScribblePrompt-Unet $0.81\pm 0.03$
ScribblePrompt-SAM $\bm{0.87\pm 0.02}$
5.2 Simulated Interactions
Setup. We use the scribble and click simulator to evaluate the performance of ScribblePrompt and baselines with iterative scribbles across the 12 evaluation sets.
For scribble-focused prompting procedures, we use:
  • Line Scribbles: Three positive line scribbles and three negative line scribbles to start, followed by one correction line scribble at each step.
  • Centerline Scribbles: One positive and one negative centerline scribble to start, followed by one correction centerline scribble at each step.
  • Contour Scribbles: One positive and one negative contour scribble to start, followed by one correction contour scribble at each step.
We limit each scribble to cover a maximum of 128 pixels. For centerline and contour scribbles, a scribble may contain multiple disconnected components.
For click-focused prompting procedures, we use:
  • Center Clicks: One positive click in the center of the largest component to start, followed by one (positive or negative) correction click per step in the center of the largest component of the error region.
  • Random Clicks: One random positive click to start, followed by one (positive or negative) correction click per step, randomly sampled from the error region.
  • Warm Start: Three random positive clicks and three random negative clicks to start, followed by one (positive or negative) correction click per step in the center of the largest component of the error region.
Since MIDeepSeg crops the input image based on the input clicks (or scribbles), it cannot make accurate predictions with only a few clicks, so we exclude it from simulations that start with a single click. For each example and interaction procedure, we simulated series of iterative interactions (10 steps for scribble-focused procedures and 20 steps for click-focused procedures) with five random seeds.
Results. As shown in Figure 4, both versions of ScribblePrompt outperform baseline methods for all simulated interaction procedures at all numbers of interactions. We show qualitative examples with simulated interactions in Figure 1 and Figure 3. We show additional results with simulated bounding boxes in Appendix D.2.
Discussion. Comparing the simulated scribble and click results (Figure 4) highlights the efficiency of scribble-based interactions. Both ScribblePrompt models reach an average Dice above $0.8$ in two scribble steps. It takes five steps of carefully-placed center clicks or seven steps of random clicks to reach the same average Dice with ScribblePrompt-SAM (Figure 18). ScribblePrompt-SAM is more robust to random clicks than ScribblePrompt-Unet, perhaps because of its pre-training with random clicks on a large number of natural images.
On average, the MIDeepSeg first prediction is poor, because the network only takes positive inputs and was trained with interior margin point clicks on a single modality. Although SAM-Med2d was trained on a large corpus of medical imaging datasets, it does not perform as well as SAM on our evaluation sets. For some tasks, such as organ segmentation, SAM-Med2d, is very accurate. However, in cases where the initial prediction is poor (e.g., retinal veins, teeth in dental X-Rays), it is less accurate than SAM or MIDeepSeg when revising its predictions based on (simulated) corrections. Detailed results by evaluation set are provided in Appendix D.3.
Figure 4: Results with simulated clicks and scribbles. We evaluate methods by simulating interactions following three scribble-focused protocols and three click-focused protocols. We show the other results, which have similar trends, in Appendix D.3. We measure Dice averaged across twelve evaluation sets, weighting each dataset equally. Shaded regions show 95% CI.
6 User Study
We conducted a user study comparing the ease of use of ScribblePrompt-Unet to SAM (ViT-b). We selected ScribblePrompt-Unet for the user study because it had similar performance and lower latency compared to ScribblePrompt-SAM. We selected SAM (ViT-b) as the baseline because it had the highest Dice in our experiments with simulated clicks and lower latency than the next closest baseline, SAM (ViT-h).
Setup. Study participants ($n=16$) were neuroimaging researchers at an academic hospital. They used each model to segment a series of nine images from nine different evaluation sets. For each segmentation task, the participants were shown the target segmentation and were asked to prompt the model until the predicted segmentation closely matched the target or they could no longer improve the prediction. We provided participants with the target segmentation to disentangle the cognitive process of identifying the region of interest from prompting the model to achieve the desired segmentation. We provide additional details and visualizations in Appendix C.
Table 2: Quantitative results of user study. We report mean Dice score and standard deviation, mean time and standard deviation (in seconds) per segmentation, and median (max) number of times the participants refreshed the prediction (per segmentation task).
SAM (ViT-b) ScribblePrompt
Dice score ($\uparrow$) $0.73\pm 0.30$ $0.84\pm 0.13$
Sec. per task ($\downarrow$) $131.7\pm 168.5$ $94.2\pm 89.0$
No. iterations ($\downarrow$) $6(47)$ $3(19)$
Results. Participants produced more accurate segmentations (0.84 vs. 0.73 Dice; $p=0.001$ using a paired t-test) using ScribblePrompt (Table 2). Participants spent $\approx 1.5$ minutes per segmentation on average using ScribblePrompt, compared to over 2 minutes per segmentation with SAM ($p=0.02$ using a paired t-test). While using ScribblePrompt, participants updated the prediction fewer times before being satisfied with the segmentation.
Upon completion, 15 out of 16 participants reported they preferred using ScribblePrompt to SAM and one participant had no preference. All participants reported it was easier to achieve the target segmentation using ScribblePrompt vs. SAM. 94.8% of participants reported ScribblePrompt was better than SAM at refining its predictions in response to scribbles, and 87.5% preferred using ScribblePrompt for clicks. 93.8% of participants reported that they felt it took less time to achieve the target segmentation using ScribblePrompt compared to using SAM.
Discussion. The most common factors that influenced participants preference for ScribblePrompt were 1) being able use multiple types of prompts including scribbles and 2) ScribblePrompt’s responsiveness to their corrections, which enabled fine-grained control over the next prediction. For some tasks, such as retinal vein segmentation, SAM was unable to make accurate predictions even with many corrections. As a result, there was more variability in segmentation accuracy and time spent per image, for SAM than ScribblePrompt.
7 Inference Runtime
Setup. We evaluate computational efficiency by measuring inference time on a single CPU for one prediction with a scribble input covering 128 pixels. Performance on a CPU reflects practical utility: because of a variety of barriers such as protected health information, users may not be able to send their data to a server for computation and must rely on local, potentially CPU-only, computing.
Results. We report performance results in Table 3. On a single CPU, ScribblePrompt-Unet requires $0.27$ seconds on average per prediction, enabling the model to be used in real-time even in low-resource environments. The next most accurate model on simulated scribbles, SAM ViT-h, requires over 2 minutes on average per prediction on a CPU. We show results with similar trends for GPU hardware in Appendix E.
ScribblePrompt-Unet achieves low latency inference because of its efficient fully-convolutional architecture. With SAM and SAM-Med2d, latency scales with the number of interactions because each point is encoded as a 256-dimensional vector embedding. For ScribblePrompt-Unet and MIDeepSeg, clicks and scribbles are encoded in masks, so inference time is constant with the number of interactions.
Table 3: Performance Summary. We measure inference time on a single CPU for a prediction with a random scribble input covering 128 pixels. We report mean and standard deviation across 1,000 runs. ScribblePrompt-SAM has the same runtime and number of parameters as SAM ViT-b.
Model Param. Runtime (sec)
SAM (ViT-h) 641M $130.79\pm 7.96$
SAM (ViT-b) 94M $13.59\pm 0.77$
SAM-Med2d w/ adapter 271M $1.23\pm 0.07$
SAM-Med2d w.o. adapter 91M $0.63\pm 0.02$
MIDeepSeg 3M $0.08\pm 0.02$
ScribblePrompt Unet 4M $0.27\pm 0.04$
8 Ablations
We conduct two ablations of ScribblePrompt: (1) varying the percentage of synthetic label inputs used during training, and (2) varying the types of prompts simulated during training. In these ablations, we report results on the validation data, specifically the validation splits of the nine validation datasets that were not used in training.
8.1 Synthetic Labels
To improve generalization to unseen tasks and datasets, we introduced synthetic labels during training. Training with some synthetic labels helps to prevent the model from over fitting, especially for small single-label datasets.
Setup. We trained ScribblePrompt-Unet with probability $p_{synth}\in\{0,0.25,0.5,1.0\}$. We evaluate models using Dice after five steps following the same six simulated interaction procedures in the main results. We show results for ScribblePrompt-SAM in Appendix B.3.2.
Results. Figure 5 shows the effect of varying the percentage of synthetic labels during training for ScribblePrompt-Unet. Using $p_{synth}=0.5$ results in the highest Dice on validation data for all simulated interaction procedures.
Figure 5: Training with some synthetic labels improves Dice. Mean change in weighted Dice relative to ScribblePrompt-Unet trained without any synthetic labels ($p_{synth}=0$). We show results for six different interaction simulation procedures at inference-time. Error bars show 95% CI. Training with both real and synthetic labels improves generalization to new datasets, compared to training with only real labels. Training with only synthetic labels results in worse Dice on new datasets compared to training with real labels.
8.2 Prompt Types
Setup. We evaluate ScribblePrompt-Unet models trained with different combinations of prompts. We consider three models trained on different subsets of prompt types:
  • ScribblePrompt-Unet (scribbles) was trained on bounding boxes and scribbles.
  • ScribblePrompt-Unet (clicks) was trained on bounding boxes and clicks following our click simulation algorithms.
  • ScribblePrompt-Unet (random clicks) was trained on bounding boxes and only random clicks.
We compare these variants of ScribblePrompt-Unet to our final ScribblePrompt-Unet model, trained on scribbles, clicks and bounding boxes.
Results. Figure 6 shows that our final version of ScribblePrompt-Unet trained with scribbles, clicks and bounding boxes predicts segmentations more accurately than do versions of ScribblePrompt-Unet trained with fewer prompt types. ScribblePrompt-Unet trained on scribbles (no clicks), has similar performance to ScribblePrompt-Unet on scribbles but worse performance on clicks. Versions of ScribblePrompt-Unet trained on our mixture of clicks or only random clicks have similar performance to ScribblePrompt-Unet on clicks.
Figure 6: Varying the types of simulated prompts during training. ScribblePrompt-Unet has similar or better performance than models trained on scribbles (no clicks) and models trained on clicks (no scribbles) for both click-focused and scribble-focused interaction scenarios. We show Dice after five steps of simulated interactions following six different (inference-time) interaction procedures. We calculate mean Dice on validation data weighting each dataset equally. Errorbars show 95% CI.
9 Conclusion
We present ScribblePrompt, a framework for interactive segmentation that enables humans to segment diverse medical images with scribbles, clicks and bounding boxes. We introduce methods for simulating realistic user interactions and generating synthetic labels. These methods enable us to train ScribblePrompt models that generalize to unseen segmentation tasks and datasets. ScribblePrompt models are more accurate than baselines, and ScribblePrompt-Unet is very computationally efficient, even on a CPU. We show in a user study that the majority of users prefer ScribblePrompt and achieve segmentations with 15% higher Dice with less perceived effort than the next most accurate baseline. ScribblePrompt promises to significantly reduce the burden of manual segmentation in medical imaging.
10 Acknowledgements
This work was supported in part by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard as well as Quanta Computer Inc. and the Wistron Corporation. Research reported in this paper was supported by the National Institute of Biomedical Imaging and Bioengineering of the National Institutes of Health under award number R01EB033773. Much of the computation resources required for this research was performed on computational hardware generously provided by the Massachusetts Life Sciences Center.
Appendix
Appendix A Demo
We showcase ScribblePrompt in a demonstration video and interactive online demo available at scribbleprompt.csail.mit.edu. The video shows ScribblePrompt-Unet being used in real time to segment test images from datasets not seen by the model during training.
Appendix B Details About ScribblePrompt Method
B.1 Prompt Simulation
In this section, we provide illustrations of the prompt simulation process. Each of these click and scribble simulation algorithms can be applied to the ground truth label (or false negative error region) to simulate positive clicks/scribbles and to the background (or false positive error region) to simulate negative clicks/scribbles.
B.1.1 Scribbles
We simulate diverse and varied scribbles by first generating clean scribbles using one of three methods: (i) line scribbles, (ii) centerline scribbles or (iii) contour scribbles. Then, we break up and warp the scribbles to add more variability to account for human error.
Line Scribbles. Figure 7 illustrates the process of simulating line scribbles.
Figure 7: Line scribbles. Given an input mask $z$, we draw random lines by sampling two end points from $\{(u,v)|z_{uv}=1\}$. We use a random deformation field to warp the line scribbles and then multiply by the binary input mask $z$ to correct parts of the scribble that were warped outside the mask. We can simulate positive scribbles by applying the algorithm to the label $y$ (top) and negative scribbles by applying the algorithm to the background $1-y$ (bottom).
Centerline Scribbles. Figure 8 illustrates the process of simulating centerline scribbles.
Figure 8: Centerline scribbles. Given an input mask, we apply a thinning algorithm [145] to get a 1-pixel wide skeleton. We break up the skeleton using a random mask and use a random deformation field to warp the broken skeleton. Lastly, we multiply the scribble mask by the input binary mask to remove parts of the scribble that were warped outside the input mask. We can simulate positive scribbles by applying the algorithm to the label $y$ (top) and negative scribbles by applying the algorithm to the background $1-y$ (bottom).
Contour Scribbles. Figure 9 illustrates the process of simulating contour scribbles.
Figure 9: Contour scribbles. We simulate a rough contour of the desired segmentation within the boundaries of the label. Given a mask $z$, we first blur the reverse mask $\bar{z}$ to reduce the size of the label such that $\tilde{z}:=1-\max(\bar{z},\bar{z}\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{z}<h$ sampled in some intensity range $h\sim U[\tilde{z}_{min},\tilde{z}_{max}]$, and use an edge detection algorithm to extract a contour inside the boundary of the given mask $z$. We break up the contour using a random mask and use a random deformation field to warp the broken contour. Lastly, we multiply the scribble mask by the input binary mask to correct parts of the scribble that were warped outside the mask. We can simulate positive scribbles by applying the algorithm to the label $y$ (bottom) and negative scribbles by applying the algorithm to the background $1-y$ (top).
B.1.2 Clicks
Interior Border Region Clicks. Figure 10 illustrates the process for simulating interior border region clicks.
Figure 10: Interior border region clicks. We sample clicks from a border region inside the boundary of a given mask. Given a mask $z$, we first blur the reverse mask $1-z$ to reduce the size of the label such that $\tilde{z}:=1-\max(1-z,(1-z)\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel with $\sigma\sim U[\sigma_{min},\sigma_{max}]$. We then sample click coordinates from $\{(u,v)|\tilde{z}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{z}_{min},\tilde{z}_{max})$. We show the simulation process for negative border region clicks on the background $1-y$ (top) and positive border region clicks on the label $y$ (bottom).
B.2 Architecture and Training
We discuss some of the modeling decisions in ScribblePrompt-Unet and ScribblePrompt-SAM.
Normalization Layers. In preliminary experiments, we evaluated normalization layers in the ScribblePrompt-Unet architecture such as Batch Norm [52], Instance Norm [124], Layer Norm [8], and Channel Norm [136]. Including normalization did not improve the mean Dice on validation data compared to using no normalization layers (Figure 11).
Figure 11: Normalization Layers. Mean Dice recorded while training ScribblePrompt-Unet with different normalization layers, averaged across five iterative predictions (using the training procedure for simulating interactions). At each epoch, we evaluate on 1,000 randomly sampled examples from the validation splits of the 65 training datasets and validation splits of the nine validation datasets. Dice was smoothed using Exponential Weighted Mean with $\alpha=0.1$.
Loss Function. In preliminary experiments, we trained ScribblePrompt with Soft Dice Loss [32], a combination of Soft Dice Loss and Binary Cross-Entropy Loss, and a combination of Soft Dice Loss and Focal Loss [72], similar to [61]. In the latter two losses, Dice Loss and BCE Loss or Focal Loss are weighted equally. We found that the combination of Soft Dice Loss and Focal Loss resulted in slightly higher mean Dice on the validation data for ScribblePrompt-Unet and ScribblePrompt-SAM. Figure 12 shows Dice recorded during training in preliminary experiments with ScribblePrompt-Unet.
Figure 12: Varying the Loss. Mean Dice recorded while training ScribblePrompt-Unet with different loss functions. We report Dice averaged across five iterative predictions (using the training procedure for simulating interactions). At each epoch, we evaluate on 1,000 randomly sampled examples from the validation splits of the 65 training datasets and validation splits of the nine validation datasets. Dice was smoothed using Exponential Weighted Mean with $\alpha=0.1$.
ScribblePrompt-SAM Details. To train ScribblePrompt-SAM, we took the pre-trained weights from the Segment Anything Model [61] with ViT-b backbone and froze all components of the network except for the decoder.
The SAM architecture can make predictions in “single-mask mode” or “multi-mask mode.” In “single-mask mode,” the decoder outputs a single predicted segmentation given an input image and user interactions. In “multi-mask mode,” the decoder predicts three possible segmentations and then outputs the segmentation with the highest predicted IoU by a MLP. We trained and evaluated ScribblePrompt-SAM in “multi-mask mode”. During training we included a MSE term in the segmentation loss to train the MLP to predict the IoU of the predictions, as in [61].
B.3 Synthetic Labels
To help reduce task overfitting – memorizing the segmentation task for single-label datasets and thus ignoring user prompts – we introduce a mechanism to generate synthetic labels. During training, for a given sample $(x^{t},y^{t})$, with probability $p_{synth}$ we replace $y^{t}$ with a synthetic label $y^{\prime}$.
B.3.1 Implementation
We use a superpixel algorithm [33] with randomly sampled scale parameter $\lambda\sim U[1,500]$ to partition the image $x^{t}$ into a map of $k$ superpixels, $z\in\{1,\dots,k\}^{n\times n}$. Then, we randomly select a superpixel as the synthetic label $y^{\prime}:=\mathbbm{1}[z=k]$. Figure 14 shows examples of training images and the corresponding maps of possible synthetic labels with different $\lambda$.
B.3.2 Synthetic Labels with ScribblePrompt-SAM
Setup. We train models with different percentages of synthetic labels during training by varying $p_{synth}$. We trained ScribblePrompt-SAM with probability $p_{synth}\in\{0,0.25,0.5\}$. We evaluate models using Dice after five steps following the same six simulated interaction procedures in the main results.
Results. Training with some synthetic labels also improves ScribblePrompt-SAM’s performance on validation data from nine (validation) datasets not seen during training (Figure 13). For ScribblePrompt-SAM, training with 50% synthetic labels leads to the highest Dice on unseen datasets at inference time.
Figure 13: Varying the amount of synthetic labels during training for ScribblePrompt-SAM. We report change in Dice relative to ScribblePrompt-SAM trained without any synthetic labels ($p_{synth}=0$). We show Dice after five steps of simulated interactions following six different (inference-time) interaction procedures. Errorbars show 95% CI.
Figure 14: Examples of possible synthetic labels. Each color in the maps is a different synthetic label. During training, we replace a given label $y$ with a synthetic label $y^{\prime}$ with probability $p_{synth}$. To generate $y^{\prime}$, we apply a superpixel algorithm with randomly sampled scale parameter $\lambda$ to the image $x^{t}$ and then randomly select a superpixel as the synthetic label. We show examples of the synthetic label maps generated using a superpixel algorithm [33] with different $\lambda$.
Appendix C User Study
Figure 15: Example segmentations and interactions from the user study. We show predictions with interactions provided by three study participants for each of the nine segmentation tasks in the user study. For each example, we visualize positive scribble and click inputs in green, negative scribble and click inputs in red, bounding box inputs in yellow, and the predicted segmentation in blue. With SAM, study participants primarily used clicks. With ScribblePrompt, participants used a mix of scribbles and clicks. For the retinal vein segmentation task, participants preferred to use clicks with both models. Participants prompted SAM with denser clicks compared to ScribblePrompt.
We conducted a user study comparing ScribblePrompt-Unet to SAM (ViT-b). We provide additional details on the design and implementation of the user study.
Study Design. The goal of the user study was to compare ScribblePrompt to the best click-focused baseline method in terms of accuracy (Dice of the final segmentations), efficiency (time to achieve the desired segmentations) and user experience (perceived effort). Participants were given time to familiarize themselves with both models on a fixed set of practice images. Afterwards they used each model to segment a series of nine test images from nine tasks (Figure 15).
The order in which participants used the models, and which image the users were assigned to segment with each model for each task was randomized. We randomly selected one training image per task to include in the set of practice images. We randomly selected two test images per task from the evaluation sets, for a total of 18 images used in the study. The models were also annonymized (i.e., “Model A” and “Model B”). We informed participants that one model was designed to be used with clicks and bounding boxes, while the other was designed for use with clicks, bounding boxes, and scribbles.
For each segmentation task, the participants were shown the target segmentation and were asked to interact with the model until the predicted segmentation closely matched the target or they could no longer improve the prediction. We provided participants with the target segmentation to disentangle the cognitive process of identifying the region of interest from prompting the model to achieve the desired segmentation.
Study Participants. Study participants were neuroimaging researchers at an academic hospital. Although the participants had prior experience with medical image segmentation, they did not necessarily have experience with the specific tasks and types of images used in the study.
We had a total of 29 participants with 16 participants completing all of the segmentations and the exit survey. We observed a higher attrition rate among participants who were assigned to use SAM first, even after being able to freely try out both models during the “practice” phase. Among the 13 participants assigned to use SAM first, 62% did not finish all of their segmentations, compared with 31% among the 16 assigned to use ScribblePrompt first. We report results on the 16 participants who completed both the segmentations and the exit survey.
Implementation. Each participant used a web-based interface powered by a Nvidia Quatro RTX8000 GPU with 4 CPUs. The interface was developed in Python using the Gradio library [2]. The interface had a “practice” mode in which users could freely switch between the two models and images from the set of practice images. After experimenting with both models, users clicked a button to begin “recorded activity” mode in which users were led through performing specific segmentation tasks with specific models. Users provided positive/negative scribble inputs, positive/negative click inputs and/or bounding box inputs, and then clicked a button to receive a prediction from the model.
Survey Results. Common factors that influenced participants preference for ScribblePrompt was being able to get accurate predictions from scribbles (“[ScribblePrompt] was more spatially smooth”), the model’s responsiveness to a variety of inputs (“it landed on my desired predictions more easily”), and less perceived effort when using the model (“[ScribblePrompt] needed much less guidance”). Participants preferred using clicks and bounding boxes over scribbles with SAM, praising its “snapiness”, the effectiveness of “exclusion clicks” and remarking it worked well for “rigid structures”. However, participants also noted in some cases “[SAM] did not respect object boundaries”, and for tasks such as retinal vein segmentation “[SAM] required lots of clicks and still was not very accurate”.
Visualizations. We visualize some of the interactions used by study participants and the resulting predictions in Figure 15. Study participants used denser clicks when prompting SAM compared to when prompting ScribblePrompt-Unet.
Appendix D Experiments
D.1 Implementation Details
Baselines. For baseline methods using the SAM architecture, we evaluate the models in both “single mask” and “multi-mask” mode. For each baseline method and interaction procedure, we selected the best performing mode based on the average Dice across the validation data and report results on test data using that mode. In the results with simulated clicks and scribbles by dataset (Figures 19, 20, 21, 22, 23, and 24), we show results using both modes.
For baseline methods using the SAM architecture, which was not designed to take scribble masks as input, we convert the scribble masks to sets of positive and negative clicks for every non-zero pixel in the scribble masks.
D.2 Simulated Bounding Boxes
We evaluate models with simulated bounding box prompts.
Setup. We evaluate segmentation accuracy using Dice score after a single bounding box prompt. We simulate bounding boxes using the same procedure as used during training ScribblePrompt: we find the minimum enclosing bounding box for the ground truth label and then enlarge each dimension by $r\sim U[0,20]$ pixels to account for human error.
In addition to SAM [61] and SAM-Med2d [25], we also include MedSAM [82] as a baseline. MedSAM is a SAM model with ViT-b backbone that was fine-tuned on a collection of over 1M medical images [^2] MedSAM used two of our evaluation datasets (TotalSegmentator and BUID) as training datasets. with one bounding box input per example. We exclude MIDeepSeg [80] from this evaluation because it is not designed to make predictions from a single bounding box input.
For methods using the SAM architecture, we apply the pixel normalization scheme in [61] to images before inference. Upon further investigation, MedSAM performed better with images rescaled to $[0,1]$; we report results for MedSAM with both normalization schemes.
Results. Figure 16 shows mean Dice after one bounding box prompt. Figure 17 shows results by dataset. ScribblePrompt-SAM has the highest Dice on average after one bounding box prompt.
Figure 16: Results with simulated bounding boxes. Mean Dice on test data from 12 datasets with one simulated bounding box prompt, weighting each dataset equally. $*{}^{*}$ indicates MedSAM with input images re-scaled to $[0,1]$ instead of the pixel normalization from [61]. Errorbars show 95% CI.
Figure 17: Results with simulated bounding boxes by dataset. Mean Dice after one simulated bounding box prompt. Among the evaluation datasets, bounding box prompts are the most effective for BUID, a breast ultrasound dataset. $*{}^{*}$ indicates MedSAM with input images re-scaled to $[0,1]$ instead of the pixel normalization from [61]. Errorbars show 95% CI.
D.3 Simulated Scribbles and Clicks
We provide more detailed results for experiments with simulated scribbles and clicks presented in the main paper.
Setup. See Section 5.2 in the main paper.
Results. Figure 18 shows Dice vs. steps of interaction for three simulated click-focused procedures and three simulated scribble-focused procedures. On average, ScribblePrompt-Unet and ScribblePrompt-SAM have the highest Dice at all steps for all of the simulated interaction procedures.
Figures 19, 20, and 21 show results by dataset for scribble-focused interaction procedures. Figures 22, 23, and 24 show results by dataset for the click-focused interaction procedures.
Figure 18: Results with simulated clicks and scribbles. We evaluate methods using three simulated scribble procedures and three simulated click procedures. We measure Dice averaged across twelve evaluation sets (the test splits of the nine validation and three test datasets), weighting each dataset equally. Shaded regions show 95% CI from bootstrapping.
Appendix E Inference Runtime
Setup. We measure inference time for a random input with a scribble covering 128 pixels. We report mean and standard deviation of inference time across 1,000 runs.
Results. We show performance results on a Nvidia RTX8000 GPU in Table 4. ScribblePrompt-Unet is faster than the baseline methods on a GPU. Unlike the transformer-based models, the inference time for each prediction is constant with the number of input interactions for ScribblePrompt-Unet.
Table 4: GPU Performance Summary. We report the mean and standard deviation for inference time on a Nvidia RTX8000 GPU for a prediction with a scribble input covering 128 pixels. ScribblePrompt SAM has the same runtime and number of parameters as SAM ViT-b.
Model Param. Runtime (ms)
SAM (ViT-h) 641M $504.36\pm 57.72$
SAM (ViT-b) 94M $133.85\pm 24.26$
MIDeepSeg 3M $65.75\pm 21.87$
SAM-Med2d w/ adapter 271M $35.06\pm 12.88$
SAM-Med2d w/o adapter 91M $24.86\pm 9.56$
ScribblePrompt-Unet 4M $1.96\pm 0.20$
Appendix F Data
We build on MegaMedical [20] to include a total of 77 datasets. We added datasets with a particular focus on Microscopy, X-Ray, and Ultrasound modalities, which were not well represented in the original MegaMedical. The full list of datasets is provided in Tables 6 and 7 and 8.
We define a 2D segmentation task as a combination of (sub)dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice for each subject. For multi-annotator datasets, we treated each annotator as a separate label. For instance segmentation datasets, we sampled one instance at a time during training. All images were resized to $128\times 128$ and re-scaled to $[0,1]$. For methods using the SAM image encoder, we convert images to RGB and apply the pixel normalization scheme in [61].
The division of data for training, validation and evaluation is summarized in 5. All 77 datasets were split into 60% train, 20% validation, and 20% test, although some splits were not used. The 77 datasets were divided into 65 training datasets (Tables 6 and 7), 9 validation datasets (ACDC, BTCV Cervix, BUID, DRIVE, HipXRay, PanDental, SCD, SpineWeb, WBC) and 3 test datasets (COBRE, SCR, TotalSegmentator). We used the “train” splits from the 65 training datasets to train the models. We use the “validation” splits from the 9 validation datasets during model development for model selection.
We report final evaluation results across 12 evaluations sets consisting of the “test” splits of the 9 validation datasets and “test” splits of the 3 test datasets to maximize the diversity of tasks and modalities in our evaluation set. No data from the 9 validation datasets or 3 test datasets were seen by the models during training.
Development Phase
Datasets Training Validation Evaluation
Training (n=65)
Validation (n=9)
Test (n=3)
Table 5: Dataset split overview. Datasets were split into 60% train, 20% validation and 20% test. Data from the “train” splits of the 65 training datasets were used to train the models. The models did not see any data from the validation or test datasets during training. Data from the “validation” split of the 9 validation datasets was used for model selection. We report final results on 12 “evaluation sets”: data from the “test” splits of the 9 validation datasets and the “test” splits of the 3 test datasets.
Table 6: Train datasets (part 1 of 2). We assembled the following set of datasets to train ScribblePrompt. For the relative size of datasets, we have included the number of unique scans (subject and modality pairs) that each dataset has.
Dataset Name Description # of Scans Image Modalities
AbdominalUS [127]
Abdominal organ segmentation
1,543
Ultrasound
AMOS [54]
Abdominal organ segmentation
240
CT, MRI
BBBC003 [77]
Mouse embryos
15
Microscopy
BBBC038 [21]
Nuclei instance segmentation
670
Microscopy
BrainDevelopment [39, 38, 64, 115]
Adult and neonatal brain atlases
53
Multimodal MRI
BrainMetShare[41]
Brain tumors
420
Multimodal MRI
BRATS [10, 11, 93]
Brain tumors
6,096
Multimodal MRI
BTCV Abdominal [66]
13 abdominal organs
30
CT
BUSIS [146]
Breast tumors
163
Ultrasound
CAMUS [67]
Four-chamber and Apical two-chamber heart
500
Ultrasound
CDemris [55]
Human left atrial wall
60
CMR
CHAOS [57, 56]
Abdominal organs (liver, kidneys, spleen)
40
CT, T2-weighted MRI
CheXplanation [112]
Chest X-Ray observations
170
X-Ray
CT2US [121]
Liver segmentation in synthetic ultrasound
4,586
Ultrasound
CT-ORG[107]
Abdominal organ segmentation (overlap with LiTS)
140
CT
DDTI [99]
Thyroid segmentation
472
Ultrasound
EOphtha [30]
Eye microaneurysms and diabetic retinopathy
102
Optical camera
FeTA [98]
Fetal brain structures
80
Fetal MRI
FetoPlac [12]
Placenta vessel
6
Fetoscopic optical camera
FLARE [81]
Abdominal organs (liver, kidney, spleen, pancreas)
361
CT
HaN-Seg [101]
Head and neck organs at risk
84
CT, T1-weighted MRI
HMC-QU [31, 60]
4-chamber (A4C) and apical 2-chamber (A2C) left wall
292
Ultrasound
I2CVB [68]
Prostate (peripheral zone, central gland)
19
T2-weighted MRI
IDRID [102]
Diabetic retinopathy
54
Optical camera
ISBI-EM [22]
Neuronal structures in electron microscopy
30
Microscopy
ISIC [26]
Demoscopic lesions
2,000
Dermatology
ISLES [46]
Ischemic stroke lesion
180
Multimodal MRI
KiTS [45]
Kidney and kidney tumor
210
CT
LGGFlair [19, 89]
TCIA lower-grade glioma brain tumor
110
MRI
LiTS [15]
Liver tumor
131
CT
LUNA [116]
Lungs
888
CT
MCIC [37]
Multi-site brain regions of schizophrenic patients
390
T1-weighted MRI
MMOTU [147]
Ovarian tumors
1,140
Ultrasound
MSD [118]
Large-scale collection of 10 medical segmentation datasets
3,225
CT, Multimodal MRI
MuscleUS [88]
Muscle segmentation (biceps and lower leg)
8,169
Ultrasound
NCI-ISBI [16]
Prostate
30
T2-weighted MRI
NerveUS [95]
Nerve segmentation
5,635
Ultrasound
OASIS [47, 86]
Brain anatomy
414
T1-weighted MRI
OCTA500 [69]
Retinal vascular
500
OCT/OCTA
PanNuke [35]
Nuclei instance segmentation
7,901
Microscopy
PAXRay [114]
92 labels covering lungs, mediastinum, bones, and sub-diaphram in Chest X-Ray
852
X-Ray
Table 7: Train datasets (part 2 of 2). We assembled the following set of datasets to train ScribblePrompt. For the relative size of datasets, we have included the number of unique scans (subject and modality pairs) that each dataset has.
Dataset Name Description # of Scans Image Modalities
PROMISE12 [74]
Prostate
37
T2-weighted MRI
PPMI [87, 29]
Brain regions of Parkinson patients
1,130
T1-weighted MRI
QUBIQ [92]
Collection of 4 multi-annotator datasets (brain, kidney, pancreas and prostate)
209
T1-weighted MRI, Multimodal MRI, CT
ROSE [83]
Retinal vessel
117
OCT/OCTA
SegTHOR [65]
Thoracic organs (heart, trachea, esophagus)
40
CT
SegThy [62]
Thyroid and neck segmentation
532
MRI, Ultrasound
ssTEM [36]
Neuron membranes, mitochondria, synapses and extracellular space
20
Microscopy
STARE [48]
Blood vessels in retinal images (multi-annotator)
20
OCT
ToothSeg [51]
Individual teeth
598
X-Ray
VerSe [78]
Individual vertebrae
55
CT
WMH [63]
White matter hyper-intensities
60
Multimodal MRI
WORD [79]
Abdominal organ segmentation
120
CT
Table 8: Validation and test datasets. We assembled the following set of datasets to evaluate ScribblePrompt and baseline methods. For the relative size of datasets, we include the number of unique scans (subject and modality pairs) that each dataset has.
Dataset Name Description # of Scans Image Modalities
ACDC [14]
Left and right ventricular endocardium
99
cine-MRI
BTCV Cervix [66]
Bladder, uterus, rectum, small bowel
30
CT
BUID [5]
Breast tumors
647
Ultrasound
COBRE [4]
Brain anatomy
258
T1-weighted MRI
DRIVE [122]
Blood vessels in retinal images
20
OCT
HipXRay [42]
Ilium and femur
140
X-Ray
PanDental [1]
Mandible and teeth
215
X-Ray
SCD [103]
Sunnybrook Cardiac Multi-Dataset Collection
100
cine-MRI
SCR [125]
Lungs, heart, and clavicles
247
X-Ray
SpineWeb [148]
Vertebrae
15
T2-weighted MRI
TotalSegmentator [134]
104 anatomic structures (27 organs, 59 bones, 10 muscles, and 8 vessels)
1,204
CT
WBC [149]
White blood cell cytoplasm and nucleus
400
Microscopy
Figure 19: Results per dataset with line scribbles. During the first step we simulate three positive line scribbles and three negative line scribbles. In subsequent steps, we simulate one (positive or negative) correction line scribble based on the error region between the previous prediction and ground truth segmentation. Each line scribble covers a maximum of 128 pixels.
Figure 20: Results per dataset with centerline scribbles. During the first step, we simulate one positive and one negative centerline scribble. In subsequent steps, we simulate one (positive or negative) correction centerline scribbles based on the error region region between the previous prediction and ground truth segmentation. Each centerline scribble covers a maximum of 128 pixels.
Figure 21: Results per dataset with contour scribbles. During the first step, we simulate one positive and one negative contour scribble based on the ground truth label. In subsequent steps, we simulate one (positive or negative) correction contour scribble based on the error region region between the previous prediction and ground truth segmentation.Each contour scribble covers a maximum of 128 pixels.
Figure 22: Results per dataset with center clicks. During the first step, one positive click is placed at the center of the largest component of the ground truth segmentation. In subsequent iterations, one (positive or negative) correction click is placed at the center of the largest component of the error region between the previous prediction and ground truth segmentation.
Figure 23: Results per dataset with random clicks. During the first step, one positive click is randomly sampled from the ground truth segmentation. In subsequent steps, one (positive or negative) correction click is randomly sampled from the error region between the previous prediction and ground truth segmentation.
Figure 24: Results with random warm start click procedure. During the first step, three positive clicks are randomly sampled from the ground truth segmentation and three negative clicks are randomly sampled from the background. In subsequent steps, one (positive or negative) correction click is placed at the center of the largest component of the error region between the previous prediction and ground truth segmentation.